# DebugAgent: Efficient and Interpretable Error Slice Discovery for Comprehensive Model Debugging

### v1

@@ A.4 More cases of the identified error slices
Figure 8: Identified slices of the image classification task by DebugAgent.
@@
Figure 10: Identified slices of the object detection task by DebugAgent.
@@
Figure 12: Identified slices of the pose estimation task by DebugAgent.
@@ A.5 Detailed experimental setup of the model repair
In the main paper, we select two model repair scenarios: data-augmented training (i.e., with the original training data) and model fine-tuning (i.e., without the original training data), across three different tasks to verify the model repair capability of DebugAgent.
For image classification, the ResNet18 model we repair is initially trained on the first 500 images per class from ImageNet (Deng et al., 2009). We utilize different image search engines to obtain a new set of 5,000 images from the web, manually ensuring no overlap with the validation set used for debugging. We then select 250 images from this set to serve as the test set, while the remaining images are designated as the query set. Both DebugAgent and HiBug, along with random selection, choose a total of 500 images (100 images per class). For pose estimation, the query and test sets are derived from an industrial private dataset, with an overall setup similar to the image classification task. For object detection, we employ RTMDet-X (Lyu et al., 2022) pretrained on the COCO dataset and fine-tuned on 1,000 images from the KITTI (Geiger et al., 2012) training set as the baseline model. We further select another 500 images from the KITTI (Geiger et al., 2012) training set for our experiment of model repair using the selection methods above and conduct a second-stage fine-tuning of the model. In this task, error slices are related to objects. For an image containing multiple objects, the average model performance regarding all objects are calculated and jointly determines the priority of the images in selection.
@@ A.6 User study
Table 6: User studies and comparisons between DebugAgent and the existing baseline HiBug
@@
We conduct a user study to evaluate the model debugging capabilities of DebugAgent in comparison to HiBug (Chen et al., 2024). The study involves 4 participants, all of whom are machine learning and computer vision practitioners. Participants were presented with the methods, a classification model, and the dataset, and they used these methods to identify error slices. We quantify the user preferences based on four criteria: (1) slice coherence, (2) slice attribute coverage, (3) slice insight, and (4) overall user experience.
@@
Results. As shown in Table 6, all of the participants prefer DebugAgent across all metrics, showing significant improvements in debugging compared to HiBug.
@@ A.7 Pseudo-code of the efficient slice enumeration algorithm and baselines
A.8 Prompts in DebugAgent
For better understanding DebugAgent, we provides three important prompts in DebugAgent. The few-shot examples are omitted. For all the prompts in DebugAgent, please refer to our code.

### v3

@@ A.4 More cases of the identified error slices
Figure 8: Identified slices of the image classification task by HiBug2.
@@
Figure 10: Identified slices of the object detection task by HiBug2.
@@
Figure 12: Identified slices of the pose estimation task by HiBug2.
@@ A.5 Detailed experimental setup of the model repair
In the main paper, we select two model repair scenarios: data-augmented training (i.e., with the original training data) and model fine-tuning (i.e., without the original training data), across three different tasks to verify the model repair capability of HiBug2.
For image classification, the ResNet18 model we repair is initially trained on the first 500 images per class from ImageNet (Deng et al., 2009). We utilize different image search engines to obtain a new set of 5,000 images from the web, manually ensuring no overlap with the validation set used for debugging. We then select 250 images from this set to serve as the test set, while the remaining images are designated as the query set. Both HiBug2 and HiBug, along with random selection, choose a total of 500 images (100 images per class). For pose estimation, the query and test sets are derived from an industrial private dataset, with an overall setup similar to the image classification task. For object detection, we employ RTMDet-X (Lyu et al., 2022) pretrained on the COCO dataset and fine-tuned on 1,000 images from the KITTI (Geiger et al., 2012) training set as the baseline model. We further select another 500 images from the KITTI (Geiger et al., 2012) training set for our experiment of model repair using the selection methods above and conduct a second-stage fine-tuning of the model. In this task, error slices are related to objects. For an image containing multiple objects, the average model performance regarding all objects are calculated and jointly determines the priority of the images in selection.
@@ A.6 User study
Table 6: User studies and comparisons between HiBug2 and the existing baseline HiBug
@@
We conduct a user study to evaluate the model debugging capabilities of HiBug2 in comparison to HiBug (Chen et al., 2024). The study involves 4 participants, all of whom are machine learning and computer vision practitioners. Participants were presented with the methods, a classification model, and the dataset, and they used these methods to identify error slices. We quantify the user preferences based on four criteria: (1) slice coherence, (2) slice attribute coverage, (3) slice insight, and (4) overall user experience.
@@
Results. As shown in Table 6, all of the participants prefer HiBug2 across all metrics, showing significant improvements in debugging compared to HiBug.
@@ A.7 Pseudo-code of the efficient slice enumeration algorithm and baselines
A.8 Prompts in HiBug2
For better understanding HiBug2, we provides three important prompts in HiBug2. The few-shot examples are omitted. For all the prompts in HiBug2, please refer to our code.